# references.bib

### v2

@@ Abstract
Remote sensing map products are used to obtain estimates of environmental quantities, such as deforested area or the effect of conservation zones on deforestation. However, the quality of map products varies, and — because maps are outputs of complex machine learning algorithms that take in a variety of remotely sensed variables as inputs — errors are difficult to characterize. Without capturing the biases that may be present, naive calculations of population-level estimates from such maps are statistically invalid. In this paper, we compare several uncertainty quantification methods — stratification, Olofsson area estimation method, and prediction-powered inference — that combine a small amount of randomly sampled ground truth data with large-scale remote sensing map products to generate statistically valid estimates. Applying these methods across four remote sensing use cases in area and regression coefficient estimation, we find that they result in estimates that are more reliable than naively using the map product as if it were 100% accurate and have lower uncertainty than using only the ground truth and ignoring the map product. Prediction-powered inference uses ground truth data to correct for bias in the map product estimate and (unlike stratification) does not require us to choose a map product before sampling. This is the first work to (1) apply prediction-powered inference to remote sensing estimation tasks, and (2) perform uncertainty quantification on remote sensing regression coefficients without assumptions on the structure of map product errors. To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a holdout ground truth dataset to be used for calibration in uncertainty quantification alongside their maps. Data and code are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
@@ 1 Introduction
However, the quality of map products varies, and resulting population-level estimates of quantities of interest may be biased and statistically invalid [jain2020benefits, bastin2017extent, venter2024uncertainty]. For example, [bastin2017extent] showed that human labeling of high-resolution imagery results in an estimate of global forest extent in dryland biomes that is 40% to 47% higher than previous estimates from remotely sensed forest cover map products. Moreover, different map products may conflict with each other and result in significantly different estimates over the same region. For example, [alix2023remotely] found that two commonly used satellite-based remote sensing maps of forest cover in Mexico differ significantly from each other, implying that there is misclassification error in at least one of the algorithms used to categorize pixels as forest or non-forest. Another recent study of three satellite-based global land cover/land use map products found that the maps are biased toward different land cover classes (e.g., one map overestimates shrub cover and another overestimates grass cover relative to the other maps) [venter2022global].
@@
Figure 1: Motivating example: map-only inference is statistically invalid. The 95% confidence interval for logistic regression coefficient relating Brazilian Amazon deforestation to slope using $N=963548$ remote sensing map points and $n=1386$ ground truth points. The imputed slope coefficient from the map alone falls outside the confidence intervals for all methods that use ground truth points.
@@
In this work, we discuss methods that correct map products with ground-truth data to give statistically-valid confidence intervals. We give a brief preview here. We seek to estimate logistic regression coefficients for the association between Brazilian Amazon deforestation in 2000–2015 and several covariates, including terrain slope. Using nearly one million remote sensing map product points and $1386$ ground truth points, we compare the $95\%$ confidence intervals for the slope coefficient with four methods (Figure 1) described shortly. See Section 4.1 for a complete description of the experimental setup.
The “classical” estimate using only ground truth points is statistically valid but has the widest confidence interval, failing to take advantage of existing map products. Meanwhile, methods that use a map product along with the ground truth points produce an estimate with lower uncertainty while still being valid.
Notably, the slope coefficient computed from the map product alone (“naive imputation”) is approximately zero, which falls outside the confidence intervals for all methods that use ground truth points. That is, if we had used only the map product, we would not have discovered the statistically significant negative relationship between slope and deforestation. Methods like PPI allow us to use ground truth points to correct for bias in the map product and thus avoid misleading conclusions.
Figure 2: Overview of uncertainty quantification using ground truth data and remote sensing map products. Olofsson area estimation, stratification, and prediction powered inference methods use a small amount of ground truth data along with a map product to produce confidence intervals for the quantity of interest. We also compare these methods against classical uncertainty quantification (using only ground truth) and naive imputation (using only the map product).
@@
In this paper, we consider uncertainty quantification methods that combine a small number of randomly sampled ground truth data ($n$ points) with large-scale remote sensing map products ($N$ points, where $N$ is much larger than $n$) in order to generate statistically valid estimates with low uncertainty (Figure 2). Stratification uses map products to determine a sampling strategy for ground truth points that reduces the variance of estimates; it has been used in the remote sensing community for decades [hay1979sampling, fitzpatrick1981comparison, stehman1998design, olofsson2014good]. Prediction-powered inference (PPI) [angelopoulos2023prediction, angelopoulos2023ppi++] is a recently developed uncertainty quantification method that uses a small set of ground truth points to calibrate and correct for bias in a much larger set of ML-generated predictions, resulting in more precise estimates than using the ground truth points alone. We compare stratification and prediction-powered inference against classical estimation (using only ground truth) and naive imputation (using only a map product).
We apply these methods to two of the most common scientific uses of remote sensing maps: area estimation and regression coefficient estimation. Area estimation includes estimating both total area occupied by a land cover/land use category and total change in land cover/land use over time. The methods applied to area estimation also apply to estimating sums or averages of continuous variables, such as total carbon stock. Regression coefficient estimation refers to estimating linear or logistic regression coefficients for causal or associative relationships between two or more environmental variables. For regressions, this paper only accounts for uncertainty in the remotely sensed dependent variable, while the covariates are assumed to have no errors. To illustrate the methods, we consider four specific remote sensing estimation tasks spanning crop area estimation, deforestation quantification, and associating deforestation and tree cover to other geophysical variables; see Table 1. Data and code for these use cases are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
@@ 2. Amazon ecoregion
Combining ground truth data with map products results in statistically valid estimates that are more reliable than naive imputation and have lower uncertainty than classical estimates. As we will show, naive imputation from maps without bias correction can result in completely incorrect area and regression coefficient estimates. To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a holdout ground truth dataset (such as the usual test set) to be used for calibration in uncertainty quantification alongside their maps, or, if that is not done, for map users to generate their own holdout set for calibration, if possible (e.g., from visual inspection of remote sensing data).
@@
Previously, Olofsson et al. [olofsson2013making] described an area estimation method that uses the confusion matrix between land cover ground truth and map product points to correct for misclassification bias in the map product. The Olofsson method is commonly used in the remote sensing community to generate confidence intervals for area estimates. However, it does not apply to regression coefficient estimation tasks. One contribution of the present work is a comparison of this method and PPI. We find both experimentally and mathematically that these two methods produce similar outputs for area estimation.
Other methods have been developed to reduce bias in regression coefficient estimates from remote sensing data. For example, [alix2023remotely] estimates the effect of a conservation program on deforestation in Mexico. They correct for bias in the remotely sensed deforestation binary variable by modelling misclassification error as dependent on environmental covariates via a logit model. Another study [gordon2023remote] estimates logistic regression coefficients for the relationship between roads and forest cover in West Africa. They train an adversarial debiasing model with ground truth data to correct for bias arising from correlations between distance from roads and the measurement error in remotely sensed forest cover. Another study [proctor2023parameter] uses co-located ground truth and remotely sensed data to show that using remote sensing data alone leads to biased regression coefficient estimates for several effect size estimation tasks in the US. They use a linear measurement error model and apply multiple imputation: a statistical method that uses a small amount of ground truth data to estimate remote sensing measurement error and impute corrected values for remotely-sensed variables. Note that this method requires an assumption about the error structure of the remotely-sensed variables, whereas the methods investigated herein do not.
Another approach to uncertainty quantification is conformal prediction, which generates statistically valid confidence sets for machine learning predictions at individual data points. For example, [valle2023quantifying] applies conformal prediction to quantify pixel-level uncertainty for a land cover/land use map product in the Brazilian Amazon, outputting a 90% predictive set of land cover classes for each pixel (i.e., the set contains the true land cover class 90% of the time). This approach differs from our work because our methods estimate total area and regression coefficients over an entire region, rather than quantifying uncertainty at individual points or pixels.
@@ 2 Background: Uncertainty Quantification
We formally describe uncertainty quantification methods that generate confidence intervals with remote sensing variables. As two baselines, we first consider methods that use only the ground-truth data (which is valid but loses information) or only use the remote-sensing data (which is not statistically valid). We then turn to methods that use both sources of data to produce statistically-valid confidence intervals. See Table 2 for a summary.
Method Data Used Statistically Valid? Caveats
@@
Stratification Ground truth points sampled
@@
Olofsson Area estimation only
@@ 2.1 Baseline methods
Naive imputation: use only the remote sensing map.
If we have a remote sensing data product with a large number $N$ of predictions $\hat{Y}_{1},\hat{Y}_{2},\dots\hat{Y}_{N}$ over the area of interest, we could treat these predictions like ground truth points and use the the formula above with $\hat{Y}$ in place of $Y$. This would produce a much smaller confidence interval, since we would use $N$ instead of $n$ in the denominator of the standard error. However, this naive imputation approach is not statistically valid, as it does not account for error or uncertainty in the model used to generate the map product predictions.
2.2 Stratification
Stratified random sampling uses remotely sensed maps to determine a sampling strategy for ground truth points that reduces the variance of estimates. Stratification is commonly used in the remote sensing community.
@@ 2.2 Stratification
The advantage of stratification with an accurate map and well-chosen sampling strategy is that the variance within each stratum will be small, which results in a smaller overall confidence interval compared to the classical method. If the map predictions are inaccurate, then the ground truth points likely have higher variance within each map stratum, resulting in a larger confidence interval. Thus the uncertainty quantification incorporates the accuracy of the map product predictions.
@@
Figure 3: Stratified random sample is a standard approach for estimation with remote sensing maps. This example shows a map of the Brazilian Amazon used for stratification. The strata are designed so that points within each stratum have low variance in deforestation status, which reduces the uncertainty when estimating area deforested. Adapted from [bullock2020satellite].
@@ 2.3 Prediction-Powered Inference
$\hat{\theta}_{\text{PPI}}=\underbrace{\frac{1}{N}\sum_{i=1}^{N}\hat{Y}_{i}^{\prime}}_{\text{naive imputation}}-\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})}_{\text{bias correction}}.$
@@
PPI can also give confidence intervals for linear and logistic regression coefficients, but the formulae are more complex. We sketch the main idea here and refer the reader to [angelopoulos2023ppi++, angelopoulos2023prediction] for a full exposition. In brief, PPI begins like naive imputation by using a large set of map predictions $\hat{Y}^{\prime}_{1},\hat{Y}^{\prime}_{2},\dots\hat{Y}^{\prime}_{N}$ to compute a measure of fit $m_{\theta}$. Then, PPI computes a bias correction $\Delta_{\theta}$ by comparing the ground truth data $Y_{i}$ to the predictions $\hat{Y}_{i}$. The relevant formula for the measure of fit and bias correction depends on what we seek to estimate — areas, binary regression coefficients, linear regression coefficients, etc. — see Table 3. Lastly, the naive $m_{\theta}$ and the bias correction $\Delta_{\theta}$ are combined to construct the PPI 95% confidence interval
@@
We use the ppi-python software package to implement the unweighted and weighted versions of PPI for our use cases. The software includes a tuning parameter $\lambda$ that we set using the map product points and ground-truth points. The confidence intervals are statistically valid even with this data-driven tuning [angelopoulos2023ppi++].
2.4 Olofsson et al. (2013) area estimation
Olofsson et al. [olofsson2013making] introduces an uncertainty quantification method that uses the confusion matrix between ground truth and map product land cover labels to estimate the area of a land cover class. This method is commonly used in the remote sensing community. Like PPI, the Olofsson method computes confidence intervals by using ground truth points to correct for bias in the map product estimates. In fact, we show mathematically in Appendix A that the two methods result in similar confidence intervals as the number of map product points $N$ grows large. However, unlike PPI, the Olofsson method only applies to area estimation but not regression coefficient estimation.
Turning to the method, suppose we have a map product with $K$ land cover classes, as well as $n$ randomly sampled ground truth points. The Olofsson method constructs the $K\times K$ misclassification confusion matrix between ground truth points and map product predictions at the ground truth point locations. For each $1\leq i\leq K$ and $1\leq j\leq K$, let $n_{ij}$ denote the number of points that are in map class $i$ and ground truth class $j$. For each class $1\leq i\leq K$, let $10<A_{i}<$ 1 denote the proportion of the total map area taken up by the map class.
Then, the area estimate for the area of class $j$ is
$\hat{\theta}_{j,\text{Olofsson}}=\sum_{i=1}^{K}A_{i}\frac{n_{ij}}{n_{i\cdot}}$

### v3

@@ Abstract
Remote sensing map products are used to obtain estimates of environmental quantities, such as deforested area or the effect of conservation zones on deforestation. However, the quality of map products varies, and — because maps are outputs of complex machine learning algorithms that take in a variety of remotely sensed variables as inputs — errors are difficult to characterize. Thus, population-level estimates from such maps may be biased. In this paper, we compare several uncertainty quantification methods — stratified estimator, post-stratified estimator, and prediction-powered inference — that combine a small amount of randomly sampled ground truth data with large-scale remote sensing map products to generate unbiased estimates. Applying these methods across four remote sensing use cases in area and regression coefficient estimation, we find that they result in estimates that are more reliable than using the map product as if it were 100% accurate and have lower uncertainty than using only the ground truth and ignoring the map product. Prediction-powered inference uses ground truth data to correct for bias in the map product estimate and (unlike stratification) does not require us to choose a map product before sampling. This is the first work to (1) apply prediction-powered inference to remote sensing estimation tasks, and (2) perform uncertainty quantification on remote sensing regression coefficients without assumptions on the structure of map product errors. To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a randomly sampled holdout ground truth dataset to be used for calibration in uncertainty quantification alongside their maps. Data and code are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
@@ 1 Introduction
However, the quality of map products varies, and resulting population-level estimates of quantities of interest may be biased [jain2020benefits, bastin2017extent, venter2024uncertainty]. For example, [bastin2017extent] showed that human labeling of high-resolution imagery results in an estimate of global forest extent in dryland biomes that is 40% to 47% higher than previous estimates from remotely sensed forest cover map products. Moreover, different map products may conflict with each other and result in significantly different estimates over the same region. For example, [alix2023remotely] found that two commonly used satellite-based remote sensing maps of forest cover in Mexico differ significantly from each other, implying that there is misclassification error in at least one of the algorithms used to categorize pixels as forest or non-forest. Another recent study of three satellite-based global land cover/land use map products found that the maps are biased toward different land cover classes (e.g., one map overestimates shrub cover and another overestimates grass cover relative to the other maps) [venter2022global].
@@
Figure 1: Motivating example: map-only estimate is biased. The 95% confidence interval for logistic regression coefficient relating Brazilian Amazon deforestation to slope using $N=963548$ remote sensing map points and $n=1386$ ground truth points. The map-only slope coefficient estimate falls outside the confidence intervals for all methods that use ground truth points.
@@
In this work, we discuss methods that correct map products with ground-truth data to give unbiased confidence intervals. We give a brief preview here. We seek to estimate logistic regression coefficients for the association between Brazilian Amazon deforestation in 2000–2015 and several covariates, including terrain slope. Using nearly one million remote sensing map product points and $1386$ ground truth points, we compare the $95\%$ confidence intervals for the slope coefficient with four methods (Figure 1) described shortly. See Section 4.1 for a complete description of the experimental setup.
The “classical” estimate using only ground truth points is unbiased but has the widest confidence interval, failing to take advantage of existing map products. Meanwhile, methods that use a map product along with the ground truth points produce an estimate with lower uncertainty while still being unbiased.
Notably, the slope coefficient computed from the map product alone is approximately zero, which falls outside the confidence intervals for all methods that use ground truth points. That is, if we had used only the map product, we would not have discovered the statistically significant negative relationship between slope and deforestation. Methods like PPI allow us to use ground truth points to correct for bias in the map product and thus avoid misleading conclusions.
Figure 2: Overview of uncertainty quantification using ground truth data and remote sensing map products. Stratified, post-stratified, and prediction powered inference estimators use a small amount of ground truth data along with a map product to produce confidence intervals for the quantity of interest. We also compare these methods against classical uncertainty quantification (using only ground truth) and map-only estimators.
@@
In this paper, we consider uncertainty quantification methods that combine a small number of randomly sampled ground truth data ($n$ points) with large-scale remote sensing map products ($N$ points, where $N$ is much larger than $n$) in order to generate unbiased estimates with low uncertainty (Figure 2). We also compare these methods against the classical estimator (using only ground truth) and map-only estimator (using only a map product). We consider the following three methods:
  • Stratified estimator uses map products to determine a sampling strategy for ground truth points that reduces the variance of estimates. This method has been used in the remote sensing community for decades [hay1979sampling, fitzpatrick1981comparison, stehman1998design, olofsson2014good].
  • Post-stratified area estimator uses the confusion matrix between ground truth and map product land cover labels to correct for misclassification bias in the map product, when estimating the area of a land cover class. This method is also commonly used in the remote sensing community [card1982using, stehman2013estimating, olofsson2013making].
  • Prediction-powered inference (PPI) [angelopoulos2023prediction, angelopoulos2023ppi++] is a recently developed uncertainty quantification method that uses a small set of ground truth points to calibrate and correct for bias in a much larger set of ML-generated predictions, resulting in more precise estimates than using the ground truth points alone.
We apply these methods to two of the most common scientific uses of remote sensing maps: area estimation and regression coefficient estimation. Area estimation includes estimating both total area occupied by a land cover/land use category and total change in land cover/land use over time. The methods applied to area estimation also apply to estimating sums or averages of continuous variables, such as total carbon stock. A map-only estimator for area is known as a pixel-counting estimator. Regression coefficient estimation refers to estimating linear or logistic regression coefficients for causal or associative relationships between two or more environmental variables. For regressions, this paper only accounts for uncertainty in the remotely sensed dependent variable, while the covariates are assumed to have no errors. To illustrate the methods, we consider four specific remote sensing estimation tasks spanning crop area estimation, deforestation quantification, and associating deforestation and tree cover to other geophysical variables; see Table 1. Data and code for these use cases are available at https://github.com/Earth-Intelligence-Lab/uncertainty-quantification.
@@ 2. Amazon ecoregion
Combining ground truth data with map products results in unbiased estimates that are more reliable than map-only estimates and have lower uncertainty than classical estimates. As we will show, map-only estimators without bias correction can result in completely incorrect area and regression coefficient estimates. To improve the utility of machine learning-generated remote sensing maps for downstream applications, we recommend that map producers provide a (simple or stratified) randomly sampled holdout ground truth dataset to be used for calibration in uncertainty quantification alongside their maps, or, if that is not done, for map users to generate their own randomly sampled holdout set for calibration, if possible (e.g., from visual inspection of remote sensing data).
@@
The remote sensing community commonly uses a post-stratified estimator to generate confidence intervals for area estimates. This estimator uses the confusion matrix between land cover ground truth and map product points to correct for misclassification bias in the map product [card1982using, stehman2013estimating, olofsson2013making]. This method only applies to area estimation, not regression coefficient estimation tasks. One contribution of the present work is a comparison of the post-stratified area estimator and PPI. We find both experimentally and mathematically that these two methods produce similar outputs for area estimation. For other methods, [gallego2004remote] provides a survey of area estimators that combine ground truth and map product data, including regression, ratio, and calibration estimators.
Methods have also been developed to reduce bias in regression coefficient estimates from remote sensing data. For example, [alix2023remotely] estimates the effect of a conservation program on deforestation in Mexico. They correct for bias in the remotely sensed deforestation binary variable by modelling misclassification error as dependent on environmental covariates via a logit model. Another study [gordon2023remote] estimates logistic regression coefficients for the relationship between roads and forest cover in West Africa. They train an adversarial debiasing model with ground truth data to correct for bias arising from correlations between distance from roads and the measurement error in remotely sensed forest cover. Another study [proctor2023parameter] uses co-located ground truth and remotely sensed data to show that using remote sensing data alone leads to biased regression coefficient estimates for several effect size estimation tasks in the US. They use a linear measurement error model and apply multiple imputation: a statistical method that uses a small amount of ground truth data to estimate remote sensing measurement error and impute corrected values for remotely-sensed variables. Note that this method requires an assumption about the error structure of the remotely-sensed variables, whereas the methods investigated herein do not.
Another approach to uncertainty quantification is conformal prediction, which generates confidence sets for machine learning predictions at individual data points. For example, [valle2023quantifying] applies conformal prediction to quantify pixel-level uncertainty for a land cover/land use map product in the Brazilian Amazon, outputting a 90% predictive set of land cover classes for each pixel (i.e., the set contains the true land cover class 90% of the time). This approach differs from our work because our methods estimate total area and regression coefficients over an entire region, rather than quantifying uncertainty at individual points or pixels.
@@ 2 Background: Uncertainty Quantification
We formally describe uncertainty quantification methods that generate confidence intervals with remote sensing variables. As two baselines, we first consider methods that use only the ground-truth data (which is unbiased but loses information) or only use the remote-sensing data (which may be biased). We then turn to methods that use both sources of data to produce unbiased confidence intervals. See Table 2 for a summary.
Method Data Used Unbiased? Caveats
@@
Stratified Ground truth points sampled
@@
Post-stratified Area estimation only
@@ 2.1 Baseline methods
Map-only estimator: use only the remote sensing map.
If we have a remote sensing data product with a large number $N$ of predictions $\hat{Y}_{1},\hat{Y}_{2},\dots\hat{Y}_{N}$ over the area of interest, we could treat these predictions like ground truth points and use the the formula above with $\hat{Y}$ in place of $Y$. This would produce a much smaller confidence interval, since we would use $N$ instead of $n$ in the denominator of the standard error. However, this map-only estimator approach may result in a biased estimate, and the confidence interval is not guaranteed to have a 95% coverage rate of the true parameter of interest. This is because the map-only estimator does not account for error or uncertainty in the model used to generate the map product predictions.
2.2 Stratified estimator
Stratified random sampling uses remotely sensed maps to determine a sampling strategy for ground truth points that reduces the variance of estimates. The stratified estimator is commonly used in the remote sensing community.
@@ 2.2 Stratified estimator
The advantage of stratified estimation with an accurate map and well-chosen sampling strategy is that the variance within each stratum will be small, which results in a smaller overall confidence interval compared to the classical method. If the map predictions are inaccurate, then the ground truth points likely have higher variance within each map stratum, resulting in a larger confidence interval. Thus the uncertainty quantification incorporates the accuracy of the map product predictions.
@@
Figure 3: Stratified estimation is a standard approach with remote sensing maps. This example shows a map of the Brazilian Amazon used for stratification. The strata are designed so that points within each stratum have low variance in deforestation status, which reduces the uncertainty when estimating area deforested. Adapted from [bullock2020satellite].
@@ 2.3 Prediction-Powered Inference
$\hat{\theta}_{\text{PPI}}=\underbrace{\frac{1}{N}\sum_{i=1}^{N}\hat{Y}_{i}^{\prime}}_{\text{map-only estimate}}-\underbrace{\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}-Y_{i})}_{\text{bias correction}}.$
@@
PPI can also give confidence intervals for linear and logistic regression coefficients, but the formulae are more complex. We sketch the main idea here and refer the reader to [angelopoulos2023ppi++, angelopoulos2023prediction] for a full exposition. In brief, PPI begins like a map-only estimator by using a large set of map predictions $\hat{Y}^{\prime}_{1},\hat{Y}^{\prime}_{2},\dots\hat{Y}^{\prime}_{N}$ to compute a measure of fit $m_{\theta}$. Then, PPI computes a bias correction $\Delta_{\theta}$ by comparing the ground truth data $Y_{i}$ to the predictions $\hat{Y}_{i}$. The relevant formula for the measure of fit and bias correction depends on what we seek to estimate — areas, binary regression coefficients, linear regression coefficients, etc. — see Table 3. Lastly, the map-only $m_{\theta}$ and the bias correction $\Delta_{\theta}$ are combined to construct the PPI 95% confidence interval
@@
We use the ppi-python software package to implement the unweighted and weighted versions of PPI for our use cases. The software includes a tuning parameter $\lambda$ that we set using the map product points and ground-truth points. The confidence intervals are unbiased even with this data-driven tuning [angelopoulos2023ppi++].
2.4 Post-stratified area estimator
The post-stratified area estimator, commonly used in the remote sensing community, utilizes the confusion matrix between ground truth and map product land cover labels to estimate the area of a land cover class [card1982using, stehman2013estimating, olofsson2013making]. Like PPI, the post-stratified area estimator computes confidence intervals by using ground truth points to correct for bias in the map product estimates. This can be viewed as a correction that uses the map errors relative to the ground truth points to modify the map-only area estimates. We show mathematically in Appendix A that PPI and the post-stratified area estimator result in similar confidence intervals as the number of map product points $N$ grows large. However, unlike PPI, the post-stratified area estimator only applies to area estimation but not regression coefficient estimation.
Turning to the method, suppose we have a map product with $K$ land cover classes, as well as $n$ randomly sampled ground truth points. The post-stratified area estimator constructs the $K\times K$ misclassification confusion matrix between ground truth points and map product predictions at the ground truth point locations. For each $1\leq i\leq K$ and $1\leq j\leq K$, let $n_{ij}$ denote the number of points that are in map class $i$ and ground truth class $j$. For each class $1\leq i\leq K$, let $10<A_{i}<$ 1 denote the proportion of the total map area taken up by the map class.
Then, the post-stratified area estimate for the area of class $j$ is
$\hat{\theta}_{j,\text{post}}=\sum_{i=1}^{K}A_{i}\frac{n_{ij}}{n_{i\cdot}}$